# Exploring System 1 and 2 communication for latent reasoning in LLMs

### v1

@@ 4.2 Reasoning evaluation: benchmarks and a controlled stress test
In this section we wanted to test if additional latent tokens improve reasoning or do they mostly add compute? We first compare our systems on GSM8K (Cobbe et al., 2021) and ProsQA (Hao et al., 2024), then use a controlled Countdown stress test (Pan et al., 2025) to probe robustness as combinatorial difficulty increases. We evaluate four systems: our reproduction of Liu et al., Hyp. 1 (frozen Base, cache-concat), Hyp. 2 (co-finetuned), and a single-model soft-embedding baseline with the same total latent budget $N_{L}$. For reference, we also compare against Coconut (GPT-2) (Hao et al., 2024). More implementation details can be found in App. 7.
Benchmarks. On GSM8K/ProsQA we follow the staged curriculum of Hao et al. (2024) and report accuracy after fine-tuning. Table 2 shows the best results ($N_{L}=12$ for this task). Figure 3A plots GSM8K accuracy versus total latents $N_{L}$; the corresponding ProsQA curves are deferred to App. 7, as this benchmark is near-saturated for competitive systems and adds limited insight.
@@
Efficiency note (Hyp. 2 vs. Coconut). Coconut generates continuous thoughts sequentially with the same network; for a given input this requires approximately $N_{L}{+}1$ full forward passes (each time appending a new latent and re-processing), so compute and latency scale linearly with $N_{L}$ Hao et al. (2024). Our Hyp. 2 uses a strict three-pass schedule independent of $N_{L}$ (Base cache $\rightarrow$ Coprocessor latents $\rightarrow$ Base decode). Assuming similar model sizes, this yields an approximate forward-pass/FLOPs reduction of
@@ 4.3 Interpretability: do latents specialize or collapse?
Reasoning can be viewed as composing distinct intermediate computations or modules (Andreas et al., 2016; Lake and Baroni, 2017; Veličković and Blundell, 2021). If our Hypothesis 2 Coprocessor truly supports such division of labor, different latents should occupy meaningfully different directions in representation space. If, instead, latents mostly scale confidence without new algorithmic structure, they will reuse the same span, yielding redundant computations (Olah et al., 2020; Elhage et al., 2022; Kornblith et al., 2019). We therefore analyze the Coprocessor’s last hidden layer (the signal fed back as input embeddings to the Base) and test whether latents specialize.
Diagnostic 1: Cross-capture heatmap. Intuition. If latents specialize, the variance of latent $j$’s activations should not lie in latent $i$’s subspace. Let $X_{i}\in\mathbb{R}^{n_{i}\times d}$ be row-centered activations for latent $i$ in this layer. For each $i$, we fit a minimal PCA subspace explaining at least $\tau=97\%$ of its variance and form the projector $P_{i}$. We then quantify how much of latent $j$’s variance falls into that subspace,
@@
Diagnostic 2: silhouette (cluster separation by latent). We compute the silhouette score (Kaufman and Rousseeuw, 1990), $s\in[-1,1]$, which for each point compares how close it is to its own latent cluster versus the nearest other latent cluster; higher $s$ means tighter, better separated clusters. This complements diagnostic 1: Instead of looking at directional variance reuse across latents (orientation overlap), it looks at instance-level spatial separation/cohesion of latent-labeled clusters. We report the global average (formal definition and the mean per-latent scores in App. §8).
@@ 4.3.1 Results
GSM8K with curriculum. Figure 4b shows the curriculum–tuned model; the no–curriculum variant is in Appendix 9. Relative to the large-scale pretraining task, fine-tuning reduces directional redundancy among latents: $\overline{H}_{\text{off}}$ drops to $0.962$ without curriculum and to $0.914$ with curriculum. However, cluster separation does not emerge: the global silhouette remains near zero in both task-tuned settings ($s\approx-0.031$). In short, curriculum nudges representations in the right direction—latents become less overlapping in orientation—but they still occupy essentially the same region of space (decorrelated directions without spatial separation).
@@ 5 Discussion
Building on Liu et al. (2024b), we cast the coprocessor architecture as a principled attempt to disentangle “System-1” token prediction from “System-2” abstract reasoning. Treating latent embeddings as a private communication channel between the two modules clarifies the conceptual link to dual-process theories in cognitive science and motivates our two new training hypotheses. However, in our setting, simply providing the channel (as in Liu et al. (2024b)) or strengthening it (our H1/H2) did not induce System-2-like computation.
@@
Our negative results do not invalidate the dual-system framework; they indicate that how the two models exchange information remains an open problem, and that current instantiations do not create conditions for System-2-like computation to emerge. Promising directions include (i) designing objectives that explicitly reward diversity or orthogonality in latent representations to encourage broader search, and (ii) developing training schedules that preserve large-scale language competence while gradually shaping latent spaces for multi-step reasoning.

### v2

@@ 4.2 Reasoning evaluation: benchmarks and a controlled stress test
In this section we wanted to test if additional latent tokens improve reasoning or do they mostly add compute? We first compare our systems on GSM8K (gsm8k) and ProsQA (coconut), then use a controlled Countdown stress test (tinyzero) to probe robustness as combinatorial difficulty increases. We evaluate four systems: our reproduction of Liu et al., Hyp. 1 (frozen Base, cache-concat), Hyp. 2 (co-finetuned), and a single-model soft-embedding baseline with the same total latent budget $N_{L}$. For reference, we also compare against Coconut (GPT-2) (coconut). More implementation details can be found in App. 7.
Benchmarks. On GSM8K/ProsQA we follow the staged curriculum of coconut and report accuracy after fine-tuning. Table 2 shows the best results ($N_{L}=12$ for this task). Figure 3A plots GSM8K accuracy versus total latents $N_{L}$; the corresponding ProsQA curves are deferred to App. 7, as this benchmark is near-saturated for competitive systems and adds limited insight.
@@
Efficiency note (Hyp. 2 vs. Coconut). Coconut generates continuous thoughts sequentially with the same network; for a given input this requires approximately $N_{L}{+}1$ full forward passes (each time appending a new latent and re-processing), so compute and latency scale linearly with $N_{L}$ coconut. Our Hyp. 2 uses a strict three-pass schedule independent of $N_{L}$ (Base cache $\rightarrow$ Coprocessor latents $\rightarrow$ Base decode). Assuming similar model sizes, this yields an approximate forward-pass/FLOPs reduction of
@@ 4.3 Interpretability: do latents specialize or collapse?
Reasoning can be viewed as composing distinct intermediate computations or modules (Andreas2016; LakeBaroni2018; Velickovic2021). If our Hypothesis 2 Coprocessor truly supports such division of labor, different latents should occupy meaningfully different directions in representation space. If, instead, latents mostly scale confidence without new algorithmic structure, they will reuse the same span, yielding redundant computations (olah2020circuits; elhage2022toy; kornblith2019similarity). We therefore analyze the Coprocessor’s last hidden layer (the signal fed back as input embeddings to the Base) and test whether latents specialize.
Diagnostic 1: Cross-capture heatmap. Intuition. If latents specialize, the variance of latent $j$’s activations should not lie in latent $i$’s subspace. Let $X_{i}\in\mathbb{R}^{N_{eval}\times d}$ be the row-centered activations for latent $i$ collected across the full evaluation set (where $N_{eval}$ is the number of task-evaluation examples). For each $i$, we fit a minimal PCA subspace explaining at least $\tau=97\%$ of its variance and form the projector $P_{i}$. We then quantify how much of latent $j$’s variance falls into that subspace,
@@
Diagnostic 2: silhouette (cluster separation by latent). We compute the silhouette score (kaufman1990finding), $s\in[-1,1]$, which for each point compares how close it is to its own latent cluster versus the nearest other latent cluster; higher $s$ means tighter, better separated clusters. This complements diagnostic 1: Instead of looking at directional variance reuse across latents (orientation overlap), it looks at instance-level spatial separation/cohesion of latent-labeled clusters. We report the global average (formal definition and the mean per-latent scores in App. §8).
@@ 4.3.1 Results
GSM8K with curriculum. Relative to the large-scale pretraining task, fine-tuning induces a noticeable reduction in directional redundancy ($\overline{H}_{\text{off}}$ drops from 0.987 to 0.914). While the global silhouette change is slim, it marks a qualitative shift, crossing from negative (0.05-0.05 without curriculum) to positive ($0.019$ with curriculum). This geometric reorganization coincides with a substantial performance gain for Hypothesis 2, improving from 11.2% (without curriculum) to 31.5%. This suggests that the curriculum induces an important shift for reasoning even if the specific dynamics of this shift remain to be fully characterized.
@@
4.4 Can Explicit Regularization Fix Latent Collapse?
To address the representational redundancy observed in §4.3 and test if lack of diversity hinders reasoning, we introduce an auxiliary orthogonality loss. We add this penalty to the standard next-token prediction objective: $\mathcal{L}_{total}=\mathcal{L}_{NTP}+\lambda\mathcal{L}_{orth}$.
Let $Z\in\mathbb{R}^{N_{L}\times d}$ be the L2-normalized latent representations. We penalize the mean squared cosine similarity between distinct latents to force them into orthogonal directions:
$\mathcal{L}_{orth}=\frac{1}{N_{L}(N_{L}-1)}\sum_{i\neq j}(Z_{i}Z_{j}^{T})^{2}$ (1)
A              B C
Refer to caption Refer to caption
Figure 5: Explicit Latent Regularization Results. A: Large scale training validation perplexity (lower is better) worsens as regularization $\lambda$ increases. B: Countdown (operands=4) accuracy restores monotonic scaling with $N_{L}$ under strong regularization ($\lambda=3$). C: Latent cross-capture heatmap for Countdown ($\lambda=3$) shows distinct, non-overlapping subspaces (compare to Fig 4).
Recovering Systematic Scaling in Reasoning (Countdown). In the combinatorial Countdown task (operands=4), explicit regularization yields a distinct phase shift. As shown in Figure 5C, a high penalty ($\lambda=3$) forces the latents into specialized, non-overlapping subspaces (Silhouette: 0.99, $\overline{H}_{\text{off}}$: 0.55). Crucially, this geometric separation translates to downstream performance: accuracy improves by 11.5 pp at $N_{L}=16$ ($62.6\%$ vs $51.1\%$). Most importantly, Figure 5B shows that regularization restores monotonic scaling with respect to the latent budget $N_{L}$. Unlike the unregularized baseline where adding latents eventually hurt performance, the regularized model effectively utilizes the additional capacity, validating that the Coprocessor can perform structured reasoning if collapse is prevented. The reader can find the evaluation accuracy score throughout training in App 12 (Figure 16).
The Trade-off with General Modeling. This specialization, however, appears antagonistic to large-scale language modeling. In large-scale pretraining (Figure 5A), increasing $\lambda$ degrades validation perplexity. This suggests that for general language distributions, the model prefers using latents as redundant "bandwidth" to propagate confident predictions (a semantic shortcut) rather than distinct reasoning steps. Consequently, the "System 2" structure required for reasoning might be locally suboptimal for pretraining, creating a tension where the objective naturally favors the "System 1" geometry of redundancy.
Interaction with Curriculum (GSM8K). On GSM8K, we found no systematic improvement ($30.4\pm 1.4\%$ across $\lambda\in[0,3]$). We hypothesize this is because the curriculum already imposes a "block-wise" structure (grouping latents by reasoning steps), which may conflict with the atomized, token-wise orthogonality enforced by $\mathcal{L}_{orth}$. Future work must explore objectives that encourage diversity between reasoning steps while allowing redundancy within a computational block.
@@ 5 Discussion
Building on liu2024deliberationlatentspacedifferentiable, we cast the coprocessor architecture as a principled attempt to disentangle “System-1” token prediction from “System-2” abstract reasoning. Treating latent embeddings as a private communication channel between the two modules clarifies the conceptual link to dual-process theories in cognitive science and motivates our two new training hypotheses. However, in our setting, simply providing the channel (as in liu2024deliberationlatentspacedifferentiable) or strengthening it (our H1/H2) did not induce System-2-like computation.
@@
Our results do not invalidate the dual-system framework; they indicate that how the two models exchange information remains an open problem, and that current instantiations do not create conditions for System-2-like computation to emerge. Our preliminary experiments on latent regularization and curriculum learning support this view, suggesting promising directions by (i) designing objectives that explicitly reward diversity or orthogonality in latent representations to encourage broader search, and (ii) developing training schedules that preserve large-scale language competence while gradually shaping latent spaces for multi-step reasoning.
[…]
6 Large scale training details
6.1 Detailed formulation of the training objective
Here we formalize the three-pass process described in Section LABEL:sec:methods. Let $x=(x_{1},\dots,x_{S})$ be a sequence. We select a set $\mathcal{T}=\{t_{1},\dots,t_{M}\}$ of $M$ augmentation indices, sampled uniformly from valid sequence positions.
(i) KV-cache generation: We compute the frozen Key-Value cache for the entire sequence once: $KV_{\theta}(x)=B_{\theta}(x)$.
(ii) Latent augmentation: For each augmentation site $t_{m}\in\mathcal{T}$, the Coprocessor generates a sequence of latents $Z^{(m)}$. It attends to the Base cache prefix $KV_{\theta}(x_{<t_{m}})$ and $N_{L}$ learnable soft tokens $S_{soft}$:
$Z^{(m)}=C_{\phi}(KV_{\theta}(x_{<t_{m}})\oplus S_{soft})\in\mathbb{R}^{N_{L}\times d}$ (2)
(iii) Decoding: The Base model predicts the next $N_{A}$ tokens (the “ahead” tokens) conditioned on the injected latents. The function $\text{inject}(\cdot)$ abstracts the specific interaction mechanism (embedding injection for Hyp. 2 or cache concatenation for Hyp. 1). The loss is computed only over these $N_{A}$ tokens for every augmentation site $m$, maximizing the likelihood:
$\mathcal{L}=\sum_{m=1}^{M}\sum_{j=1}^{N_{A}}\log p_{\theta}\left(x_{t_{m}+j}\mid\text{inject}\left(KV_{\theta}(x_{<t_{m}}),Z^{(m)}\right),x_{t_{m}:t_{m}+j-1}\right)$ (3)
$M$ allows us to train on multiple randomly sampled reasoning segments per sequence in parallel, while $N_{A}$ defines the window of future tokens that provide the supervisory signal.
6.2 Implementation details
Figure 6: Soft-embedding baseline (main control). A single transformer (no Coprocessor) plays both “System-1” and “System-2” roles. We attach $N_{L}$ learnable soft tokens (“latent slots”) to the Base and continue pre-training the same Base checkpoint LLM and these tokens on the same dataset and token budget as the dual-model runs. This matches the dual setup’s latent capacity ($N_{L}$) while using half the trainable parameters, directly testing whether a separate coprocessor—and its cache-level latent communication with the Base—provides benefits beyond added compute. During decoding, the learned soft tokens serve as the latent context via input-embedding injection.
Figure 7: Three-pass schedule on a toy sequence. Example with input “abcde” and two augmentation sites at b and d (M=2M{=}2), $N_{L}{=}2$ latent slots per site ($\texttt{b}^{\prime},\texttt{b}^{\prime\prime}$, $\texttt{d}^{\prime},\texttt{d}^{\prime\prime}$), and $N_{A}{=}1$ ahead token per site. Pass 1 (Base): run once on the raw sequence to form the cache $KV_{\theta}(x)$. Pass 2 (Coprocessor): consume $KV_{\theta}(x)$ together with the concatenated $M\!\times\!N_{L}$ latent placeholders $\texttt{b}^{\prime},\,\texttt{b}^{\prime\prime},\,\texttt{d}^{\prime},\,\texttt{d}^{\prime\prime}]$, producing per-slot latent embeddings $z_{\texttt{b}^{\prime}},z_{\texttt{b}^{\prime\prime}},z_{\texttt{d}^{\prime}},z_{\texttt{d}^{\prime\prime}}$. Pass 3 (Base decode): reuse the same $KV_{\theta}(x)$ and feed the concatenated outputs plus all $M\!\times\!N_{A}$ ahead tokens (one per site in this toy) as input embeddings (e.g., $z_{\texttt{b}^{\prime}},z_{\texttt{b}^{\prime\prime}},z_{\texttt{d}^{\prime}},z_{\texttt{d}^{\prime\prime}},\hat{\texttt{c}},\hat{\texttt{e}}$]). Concatenation in Passes 2–3 enables many augmentations per sequence to be trained in parallel while keeping latent computation (Pass 2) separated from next-token prediction (Pass 3). We are still unsure by the single-pass implementation sketched by liu2024deliberationlatentspacedifferentiable (no public code), but this schedule preserves disentanglement without sacrificing batch parallelism.
7 Reasoning implementation details
Benchmarks and curriculum.
For GSM8K gsm8k and ProsQA coconut we follow the staged curriculum of coconut. At stage $k$, the first $k$ CoT steps are replaced by $k\times c$ continuous latents. Loss is masked on question tokens and on the latent spans; only the remaining language tokens contribute to the objective. We sweep $c\!\in\!\{1,2,3,4\}$ with k=4k{=}4, so the total latent budget is $N_{L}{=}k\,c\!\in\!\{4,8,12,16\}$. All benchmark fine-tuning runs use a learning rate of $1{\times}10^{-4}$. To make the comparison to coconut as faithful as possible, we deliberately minimize changes to their curriculum recipe; better schedules may exist for our dual-architecture, but exploring them is out of scope here.
Figure 8: ProsQA accuracy vs. total latents $N_{L}$.
Notes. ProsQA is near-saturated for both GPT-2 and Qwen-3, so curves are flat across $N_{L}$ and the soft-embedding baseline tracks Hypothesis 2 closely. This complements Fig. 3A (GSM8K), where accuracy likewise fails to grow monotonically with $N_{L}$. Plotting conventions match the main text; dotted lines denote the single-model soft-embedding baseline.
Countdown setup.
Countdown is trained without curriculum. Operands are sampled uniformly from $[1,50]$ and the target from $[0,100]$; expressions use ++, -, $\div$ and $\times$ (as in TinyZero tinyzero). For each $(\text{operands},N_{L})$ setting we train for $3$ epochs over 262,144262{,}144 generated training samples and evaluate on 1,0241{,}024 held-out samples. We vary the operand count $\in\{3,4,5\}$ and latent budget $N_{L}\!\in\!\{1,2,4,8\}$. All runs use a learning rate of $1{\times}10^{-4}$. The branching factor grows with operands as $\mathrm{Catalan}(n{-}1)\,2^{n-1}\,n!$, yielding the controlled difficulty axis reported in Fig. 3B.
Fine-tuning from large-scale pre-training.
On GSM8K/ProsQA, resuming from large-scale pre-training hurts relative to starting from base checkpoints (Tables 34), suggesting a mismatch between next-token pre-training geometry and our curriculum-based supervision at this scale.
Comparison protocols (all end curriculum at $N_{L}{=}16$).
  • Hyp. 2 (scratch): co-finetune Base and Coprocessor from the standard base checkpoints; staged curriculum with k=4k{=}4 stages and c=4c{=}4 latents per stage.
  • Resume + curriculum: initialize Hyp. 2 from the large-scale training checkpoints (Sec. 4.1) and run the same curriculum (k=4k{=}4, c=4c{=}4).
  • Straight-to-16: initialize from the large-scale checkpoints but skip the curriculum; a single stage with k=1k{=}1, c=16c{=}16.
Table 3: Effect of resuming from large-scale pre-training on GSM8K. Accuracy (%); $\Delta$ is the change vs. Hyp. 2 (scratch).
Model Hyp. 2 (scratch) Resume + curriculum Straight-to-16
GPT-2 31.5 26.0 ($\Delta\!-5.5$) 28.2 ($\Delta\!-3.3$)
Qwen-3 38.6 35.7 ($\Delta\!-2.9$) 37.3 ($\Delta\!-1.3$)
Table 4: Effect of resuming from large-scale pre-training on ProsQA. Accuracy (%); $\Delta$ is the change vs. Hyp. 2 (scratch).
Model Hyp. 2 (scratch) Resume + curriculum Straight-to-16
GPT-2 99.0 97.1 ($\Delta\!-1.9$) 97.3 ($\Delta\!-1.7$)
Qwen-3 99.5 88.3 ($\Delta\!-11.2$) 84.5 ($\Delta\!-15.0$)
By contrast, on Countdown (no curriculum) resuming helps at $N_{L}{=}16$: Fig. 9 compares three curves per family (GPT-2 left, Qwen-3 right)—Hypothesis 2 (from scratch), Hypothesis 1, and Hypothesis 2 resumed from large-scale training—and the resumed Hypothesis 2 dominates. One plausible explanation is that Countdown’s objective is closer to the pre-training signal (no curriculum, homogeneous task family), whereas GSM8K/ProsQA curriculum alters the supervision structure in a way that conflicts with the pre-trained latent geometry. A fuller exploration of how to transfer large-scale pre-training into reasoning-specific curricula is a promising direction for future work.
Figure 9: Countdown appendix: accuracy at $N_{L}{=}16$ for Hypothesis 2 (from scratch), Hypothesis 1, and Hypothesis 2 resumed from large-scale pre-training. Left: GPT-2. Right: Qwen-3. Resuming improves Countdown despite hurting GSM8K/ProsQA, hinting at a mismatch between curriculum-based supervision and pre-trained latent geometry.
8 Silhouette definition
Let $X=\{x_{p}\}_{p=1}^{N}\subset\mathbb{R}^{d}$ be all occurrences from the coprocessor’s last hidden layer, labeled by latent indices $y_{p}\in\{1,\dots,N_{L}\}$ with clusters $C_{\ell}=\{p:\,y_{p}=\ell\}$. For $p\in C_{\ell}$, define the mean intra-latent distance
$a_{p}\;=\;\frac{1}{|C_{\ell}|-1}\sum_{\begin{subarray}{c}q\in C_{\ell}\\ q\neq p\end{subarray}}\!\!\|x_{p}-x_{q}\|_{2},$
and, for $\ell^{\prime}\neq\ell$, the mean distance
$d(p,C_{\ell^{\prime}})\;=\;\frac{1}{|C_{\ell^{\prime}}|}\sum_{q\in C_{\ell^{\prime}}}\|x_{p}-x_{q}\|_{2},\qquad b_{p}\;=\;\min_{\ell^{\prime}\neq\ell}d(p,C_{\ell^{\prime}}).$
The silhouette of $p$ is
$s_{p}\;=\;\frac{b_{p}-a_{p}}{\max\{a_{p},\,b_{p}\}}\in[-1,1].$
We report the global silhouette $s=\tfrac{1}{N}\sum_{p=1}^{N}s_{p}$ and the per-latent silhouette $s_{\ell}=\tfrac{1}{|C_{\ell}|}\sum_{p\in C_{\ell}}s_{p}$. For singleton clusters, we set $s_{p}=0$.
9 Additional interpretability results
9.1 Effective Subspace Dimensions
To verify that the high cross-capture values ($\bar{H}_{off}$) observed in Figure 4 are not artifacts of high-dimensional subspaces trivially covering the ambient space (where $P_{i}\approx I$), we analyze the dimensionality of the PCA subspaces used in our metric.
For each latent $i$, we compute the rank $k$ of the projector $P_{i}$ required to explain $\tau=97\%$ of the variance. Table 5 reports the average rank across all latents for the GPT-2 backbone ($d_{model}=768$).
Table 5: Average dimensionality of latent PCA subspaces (explaining 97% variance) for GPT-2 ($d=768$).
Task Avg. Subspace Rank % of Ambient Dim
Large Scale Pretraining $\approx 2$ $0.2\%$
Countdown $\approx 60$ $7.8\%$
GSM8K (Curriculum) $\approx 161$ $20.9\%$
In all settings, the effective subspace dimension is significantly smaller than the ambient dimension. This confirms that the high off-diagonal capture we observe implies specific, redundant directional alignment between latents, rather than broad spatial coverage.
9.2 Main paper tasks
Figure 10: Supplementary latent cross-subspace capture heatmaps (coprocessor last layer). A: Countdown with operands = 3. B: Countdown with operands = 5. C: GSM8K without curriculum. Panels A/B are qualitatively similar to the main-paper operands = 4 plot; panel C shows stronger collapse than its curriculum counterpart.
Latent Large-scale pretrain GSM8K (curr.) Countdown (op=4)
1 -0.158 -0.014 0.912
2 -0.089 -0.036 0.506
3 -0.176 -0.036 0.324
4 -0.173 -0.032 0.209
5 -0.259 -0.044 0.243
6 -0.113 -0.047 0.550
7 -0.067 -0.047 0.635
8 -0.143 -0.045 0.629
9 -0.117 -0.032 0.723
10 -0.170 -0.040 0.742
11 -0.202 -0.022 0.568
12 -0.252 -0.019 0.357
13 -0.193 0.336
14 -0.133 0.461
15 -0.191 0.026
16 -0.211 0.030
Table 6: Per-latent silhouette $s_{\ell}$ . Dashes indicate unused latents in the GSM8K run (here $N_{L}{=}12$). Large-scale pretraining shows uniformly negative silhouettes (cluster intermixing); GSM8K with curriculum hovers near zero (weak separation); Countdown exhibits strong separation for most latents while still showing redundancy in cross-subspace capture (see main-text heatmaps).
9.3 Extra Reasoning Tasks: SVAMP and MultiArith
Setup. We fine-tuned the Qwen-3 dual-model variant (Hypothesis 2) on the training sets of these tasks. Unlike the GSM8K experiments in the main text (which utilized a staged curriculum derived from CoT data), these models were fine-tuned directly on the target data without a curriculum. We report results exclusively for Qwen-3; we found that the smaller GPT-2 model failed to improve over random baselines on these datasets. We attribute this to the limited number of training examples available in SVAMP and MultiArith compared to Countdown, and the absence of a CoT-based curriculum which was essential for stabilizing optimization on GSM8K.
Figure 11: Accuracy scaling with respect to latent count $N_{L}$ on SVAMP and MultiArith (Qwen-3) after finetuning on 3 training epochs. Similar to GSM8K and Countdown, we observe no systematic benefit from increasing the latent budget; performance either plateaus (SVAMP) or degrades (MultiArith), consistent with the hypothesis that the latents are suffering from representational collapse.
Results: Latent Collapse. Despite learning the tasks to a moderate degree, the latent space analysis at $N_{L}=16$ reveals the same pattern of collapse observed in the main paper:
  • MultiArith: The latents exhibit high redundancy, with a mean off-diagonal capture of $\overline{H}_{off}=0.950$ and a global silhouette score of s=0.0114s=-0.0114.
  • SVAMP: Similarly, we observed a mean off-diagonal capture of $\overline{H}_{off}=0.949$ and a global silhouette score of s=0.056s=-0.056.
Figure 12: Latent cross-subspace capture heatmaps for A) SVAMP and B) MultiArith (Qwen-3). Both tasks display high off-diagonal brightness ($\overline{H}_{off}\approx 0.95$), indicating that the "latent collapse" phenomenon generalizes to other arithmetic reasoning benchmarks.
To address the question of whether the observed representational redundancy is specific to our main datasets or a general property of the architecture, we extended our interpretability analysis to two additional arithmetic reasoning benchmarks: SVAMP and MultiArith.
Results: Latent Scaling. We swept the latent budget $N_{L}\in\{2,4,8,16\}$ to test if additional latent capacity translates to better reasoning. As shown in Figure 11, we observe the same "flat or degrading" scaling law seen in the main text:
  • SVAMP: Performance remains largely stagnant as $N_{L}$ increases, peaking slightly at $N_{L}=4$ ($\approx 45\%$) before flattening out around $42\%$.
  • MultiArith: Accuracy actually degrades as latent capacity increases, dropping from $\approx 28\%$ at $N_{L}=2$ to $20\%$ at $N_{L}=16$.
This confirms that without explicit regularization or curriculum to shape the latent space, simply adding more "thinking tokens" does not yield robust performance gains.
10 GPT-2 with matched training parameters
To strictly verify whether the dual-model architecture offers benefits beyond simply increasing the parameter count, we compare our 124M dual-model setup against a single, larger GPT-2 model.
Setup. We train a “small to medium” GPT-2 with 16 layers, 16 heads, and 1024-dimensional embeddings, resulting in approximately 254M parameters. This matches the aggregate parameter count of our dual-system (124M Base + 124M Coprocessor). We first pre-train this model from scratch using the NanoGPT framework (karpathy2022nanogpt) to serve as a parameter-matched baseline. We note that we limited this ablation to the GPT-2 family, as performing full pre-training of a parameter-matched Qwen model ($\approx$ 1.2B parameters) from scratch is unrealistic for the scope of this paper.
Results. As shown in Figure 13, the parameter-matched single model (Soft embeddings 250M) attains lower validation perplexity than the Hypothesis 2 dual-system.
A Conservative Comparison. We acknowledge that this comparison places the dual-model system at an initialization disadvantage. While the Base component is pre-trained, the Coprocessor (comprising 50% of the aggregate parameters) is initialized randomly and must learn from scratch. In contrast, the single 250M model is trained as a fully coherent unit. This is our best attempt at having a matched parameter baseline.
Figure 13: Validation perplexity for GPT-2 variants vs. a parameter-matched single model. We compare the 124M dual-model runs against a single 250M GPT-2 (dotted lines). The single model attains lower perplexity, suggesting that the dual architecture is an inefficient way to scale parameters compared to a unified model of the same size.
11 Extra ablations
A                    B
Figure 14: Validation perplexity whilst training on the FinWeb-Edu-100BT corpus. A: Ablating number of latents $N_{L}$ of the GPT-2 Coprocessor for Hypothesis 1. B: Ablating number of latents $N_{L}$ of the Qwen-3 Coprocessor for Hypothesis 1.
Model Cont. pretraining $N_{A}=32$ $N_{A}=16$ $N_{A}=8$ $N_{A}=4$
GPT-2 30.90 31.29 31.25 31.04 30.93
(+0.39) (+0.35) (+0.14) (+0.03)
Qwen-3 34.50 52.20 53.10 48.20 46.80
(+17.70) (+18.60) (+13.70) (+12.30)
Table 7: ARC-Easy accuracy (in %) for varying numbers of ahead tokens $N_{A}$ during coprocessor training. For GPT-2, $N_{A}=32$ achieves the highest accuracy (31.29, +0.39 points over the Cont. pretraining baseline of 30.90). For Qwen-3, $N_{A}=16$ achieves the highest accuracy (53.10, +18.60 points over the Cont. pretraining baseline of 34.50).
11.1 Length Generalization Analysis
To address concerns regarding the robustness of our proposed architectures—specifically whether the models are sensitive to sequence length scaling—we evaluate our trained models on sequence lengths extending beyond the training context.
All models were trained with a sequence length of $S=512$ despite the model supporting up to 1024. This design choice ensures that our generalization tests (up to $2\times S=1024$) isolate the effect of our training method from the absolute positional limits of the pre-trained checkpoint. We evaluate validation perplexity on the FineWeb-Edu corpus with sequence lengths scaled by factors of $\{1.0\times,1.125\times,1.25\times,1.5\times,2.0\times\}$ (evaluated up to 1024 tokens).
Figure 15: Length generalization tests evaluating validation perplexity on sequence lengths up to $2\times$ the training context ($S=512$). Top Row: GPT-2 variants. Bottom Row: Qwen-3 variants. Left Column: Hypothesis 1 (Frozen Base + KV Concat). Right Column: Hypothesis 2 (Co-finetuning).
The results are summarized in Figure 15:
  • Hypothesis 1 (Frozen-Base KV Augmentation): Contrary to concerns that cache concatenation might disrupt attention mechanisms at longer lengths, Hypothesis 1 demonstrates strong robustness. For Qwen-3, performance remains stable across all tested lengths. For GPT-2, the model generalizes well up to $1.5\times$ length, with only slight degradation observed at $2\times$ length.
  • Hypothesis 2 (Co-finetuned): The fully finetuned dual-model system shows architecture-dependent generalization. On Qwen-3 (which uses Rotary Positional Embeddings), H2 generalizes effectively. However, on GPT-2 (which uses absolute positional embeddings), H2 fails to generalize, exhibiting rapid perplexity degradation when inputs exceed the training context.
This suggests that these dual-model architectures are robust to length generalization provided they utilize modern relative positional embeddings (like RoPE in Qwen-3), whereas architectures with absolute positional embeddings (as in GPT-2) may struggle to generalize in the co-finetuned setting (Hyp 2).
12 Training Dynamics of Regularization
To understand whether the orthogonality loss impacts convergence speed or stability, we track the evaluation accuracy on the Countdown task (operands=4) throughout the training process. Figure 16 illustrates the learning curves for varying regularization strengths $\lambda\in\{0,0.03,0.3,1,3\}$.
We observe that while strong regularization ($\lambda=3$) initially penalizes performance, leading to a slower start compared to the unregularized baseline ($\lambda=0$), it prevents the early saturation seen in the baseline. The unregularized model converges rapidly to a lower performance ceiling, likely settling into a redundant local minimum. In contrast, the regularized models continue to improve later into training.
Figure 16: Countdown evaluation accuracy (Operands=4 for $N_{L}=16$) tracked over training steps for different orthogonality regularization strengths $\lambda$. While the baseline ($\lambda=0$) converges quickly, stronger regularization ($\lambda=3$) sustains improvement for longer, ultimately reaching higher accuracy.